# A Dynamic, Self-balancing k-d Tree

### v13

@@ 1 Introduction
Bentley introduced the k-d tree as a binary search tree that stores k-dimensional data [Bentley]. Like a binary search tree, a k-d tree partitions a set of data at each recursive level of the tree. Unlike a binary search tree that uses only one key for all levels of the tree, a k-d tree uses $k$ keys and cycles through the keys for successive levels of the tree. For example, to build a k-d tree from three-dimensional tuples that represent $\left(x,y,z\right)$ coordinates, the keys would be cycled as $x,y,z,x,y,z$... for successive levels of the k-d tree. A more elaborate scheme for cycling the keys chooses the coordinate that has the widest dispersion or largest variance to be the key for a particular level of recursion [Friedman] but this article assumes that the keys are cycled as $x,y,z,x,y,z$... in order to simplify the exposition.
@@
Two dynamic k-d trees proposed previously maintain balance via a set of multiple k-d trees [Procopiuc] or k-d* trees [Willard]. In contrast, the algorithms described in the present article maintain balance via a single k-d tree.
@@ 1.1 Two Definitions of Balance
This definition of height permits definition of two types of balance. The first type of balance, AVL balance, requires that each node’s left and right subtrees differ in height by at most one node [Adelson] [Willard].
The second type of balance, red-black balance, requires that, at each node of the tree, the node’s left and right subtrees differ in height by a factor of 2 or less. This requirement derives from the following features of a red-black tree: (1) a node is colored either red or black; (2) if a node is red, neither of its children may be red; and (3) every path from a given node to the bottom of the tree must contain the same number of black nodes [Guibas]. Given these features, the shortest path from any node to the bottom of a red-black tree contains only black nodes, whereas the longest path contains alternating red and black nodes. Hence, because every path from any node to the bottom of the tree must contain the same number of black nodes, the longest path is twice the length of the shortest path [Drozdek].
@@ 1.2 Insertion into a k-d Tree and Subsequent Rebalancing
Insertion into a k-d tree occurs in a similar manner to insertion into a standard binary search tree (i.e., a 1-d binary search tree) [Drozdek] [Weiss]. To insert a k-dimensional tuple, the tree is searched recursively for that tuple, comparing the tuple’s super key to the tree node’s $x$:$y$:$z$, $y$:$z$:$x$, or $z$:$x$:$y$ super key at each level of the tree as the tuple descends the tree. If the search arrives at the bottom of the tree without finding the tuple, the tuple is inserted into the tree as a new leaf node.
Then, as the recursion unwinds, the height is computed at each node along the return path to the root of the tree, and the balance is checked at each node to determine whether rebalancing is required at that node. In Figure 1, the difference in child subtree heights is 0, 1, and 2 at nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ respectively. Hence, the subtree rooted at node $\left(8,1,5\right)$ must be rebalanced, as required by the AVL or red-black criterion because either criterion enforces a subtree height difference of at most one node. This subtree is rebalanced by rebuilding it [Overmars] to create the subtree shown in Figure 2 that contains nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ depicted with dashed borders, and whose respective $z$:$x$:$y$ super keys have relative order 11:99:22 << 22:88:33 << 55:88:11. Rebuilding a subtree that contains 3 nodes or fewer, such as this subtree, occurs frequently and requires comparison of at most 3 super keys [Stepanov]. Rebuilding a subtree that contains more than 3 nodes requires a comprehensive tree-building algorithm, such as a $O\left[kn\log\left(n\right)\right]$ or $O\left[n\log\left(n\right)\right]$ parallel algorithm [Brown2015].
@@ 1.3 Deletion from a k-d Tree and Subsequent Rebalancing
Deletion from a k-d tree begins in a similar manner to, but is more complicated than, deletion from a 1-d binary search tree [Drozdek], [Rocca] [Weiss]. To delete a k-dimensional tuple, the tree is searched recursively for that tuple, comparing the tuple’s super key to the node’s $x$:$y$:$z$, $y$:$z$:$x$, or $z$:$x$:$y$ super key at each level of the tree. If the search finds the tuple, the tuple is deleted from the tree as one of three cases: (1) a leaf node with no children, (2) a node with one child, and (3) a node with two children. After deletion of the node, the height is computed at each node along the return path to the root of the tree as the recursion unwinds, and the balance is checked at each node to determine whether rebalancing is required at that node.
@@
The second case is deletion of a node that has one child. This case must be treated differently for a k-d tree than for a 1-d tree. In a 1-d tree, a 1-child node is replaced by its only child. However, for a k-d tree, replacing a 1-child node by its child violates the invariant of the subtree rooted at that child [Drozdek] [Rocca]. A 1-child node, whose super key is $x$:$y$:$z$, has a child whose super key is $y$:$z$:$x$. So moving that child to the next higher level in the tree to replace its parent, as is done in a 1-d tree, would require that the child’s super key be $x$:$y$:$z$ instead of $y$:$z$:$x$.
This problem is solved by replacing a 1-child node by either its immediate predecessor node or its immediate successor node, similar to the replacement of a node that has two children in a 1-d tree. A 1-child node that has only a less-than child is replaced by its immediate predecessor, whereas a 1-child node that has only a greater-than child is replaced by its immediate successor [Drozdek] [Rocca].
However, finding the predecessor or successor node is more complicated for a k-d tree than for a 1-d tree [Drozdek]. The predecessor node to a particular node in a 1-d tree is found by starting with the less-than child to that node, and recursively following the greater-than child pointer until it is null, because the null-pointer node has the largest key in the less-than subtree [Wirth]. The successor node to a particular node is found in a similar manner by starting with the greater-than child to that node, and recursively following the less-than child pointer until it is null.
@@
Because node $b$ has a $y$:$z$:$x$ super key, its $x$:$y$:$z$ sorted order relative to node $a$ is unknown, so it must inspected as a potential predecessor node to node $a$. Also, because of its $y$:$z$:$x$ super key, node $b$ cannot eliminate its children (nodes $c$ and $d$) from recursive search for node $a$’s predecessor node.
@@
In view of the exclusion of less-than branches, finding 1-child node $a$’s immediate predecessor node requires inspecting $n^{1-1/k}$ nodes of the $n$-node subtree rooted at node $b$, assuming that subtree to be perfectly balanced with no empty branches [Samet]. And if 1-child node $a$ has only a greater-than child instead of a less-than child, finding node $a$’s immediate successor node requires searching the subtree rooted at that greater-than child to find the node that has the smallest $x$:$y$:$z$ super key.
When the immediate predecessor or immediate successor node to the 1-child node has been found in the manner described above, that predecessor or successor node’s tuple field is copied to the 1-child node’s tuple field, and then the predecessor or successor node is deleted from the tree. But that predecessor or successor node is not necessarily a leaf node, and if not, it must be deleted recursively as one of the three cases itemized above [Rocca], [Samet]. Recursive deletion converges to a leaf node that is excised from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path.
The third case is deletion of a node that has two children. This case differs from the 1-child case only in that a deleted 2-child node may be replaced by either its immediate predecessor or its immediate successor node. The choice between the predecessor or successor is arbitrary. However, the optimum replacement node might be selected from the 2-child node’s child subtree that has the greater height, in an attempt to improve balance and thereby minimize rebalancing the subtree rooted at the 2-child node [Foster1965].
@@ 1.4 Benchmark Methodology
A k-d tree benchmark measured algorithm performance for trees that contained $n$ nodes, for values of $n$ in the range $\left[1,003,201;4,523,071\right]$ that map to equally spaced values of $n\log_{2}\left(n\right)$ in the range $\left[20,000,000;100,000,000\right]$. Each node of the tree stored a $k$-dimensional tuple of 64-bit integers. The integers were equally spaced across the maximum 64-bit integer range $r$, so the spacing was $r/n$. The integers were randomly shuffled via the std::mt19937_64 Mersenne Twister pseudo-random number generator [Matsumoto] and copied to the first of the $k$ dimensions, then randomly shuffled again and copied to the second of the $k$ dimensions, et cetera. So that all benchmarks randomly shuffled the integers in an identical sequence, each benchmark initialized std::mt19937_64 to std::mt19937_64::default_seed.
@@
Each benchmark measured the execution times for insertion of all tuples into the tree, verification of correct ordering of the tree, search for all tuples in the tree, and deletion of all tuples from the tree via the std::chrono::steady_clock::now() function. To provide a comparison, the execution time was measured for construction of a static k-d tree via a $O\left[n\log\left(n\right)\right]$ algorithm. Each benchmark was repeated 100 times and the mean values and standard deviations of the execution times were calculated. The standard deviations were less than 1% of the mean values.
To assess the the performance of the dynamic k-d tree for worst-case data, a set of tuples was created by constructing a static k-d tree, and then sweeping that tree in sorted order to obtain a set of tuples to insert into and delete from a dynamic k-d tree.
@@ 1.5 Benchmark Results
In Figure 4, the execution-time plots for insertion, deletion, and search of random and sorted tuples in a dynamic k-d tree, as well as for creation of a static k-d tree from random tuples, demonstrate $O\left[n\log\left(n\right)\right]$ computational complexity, as expected for a balanced binary tree. The execution time for insertion of random tuples is roughly 1.5 times longer than the execution time for creating a static k-d tree en masse from random tuples. This disparity represents the difference between rebalancing piecemeal for a dynamic tree, versus balancing once en masse for a static tree.
Also, in Figure 4, the execution time for insertion of random tuples exceeds the execution time for insertion of sorted tuples. This difference in execution times is likely due to cache memory effects. For insertion of sorted 3-dimensional $\left(x,y,z\right)$ tuples, two consecutive tuples descend the nascent tree via a similar path. Hence, the nodes that the first tuple causes to be loaded into cache during its descent of the tree likely remain resident in cache and are likely accessed by the second tuple during its descent of the tree. In contrast, for insertion of random tuples, it is unlikely that two consecutive tuples descend the tree via a similar path. Hence, the nodes loaded into cache during the first tuple’s descent of the tree are unlikely to be accessed by the second tuple. A similar disparity in execution times between sorted and random keys has been observed for the AVL and red-black trees [Brown2025].
In addition, in Figure 4, the execution time for deletion of sorted tuples exceeds the execution time for deletion of random tuples. This difference in execution times is likely due to the fact that deletion of sorted tuples requires rebuilding larger subtrees than does deletion of random tuples. Table 1 reports the size (in units of kilo-nodes) of the largest subtree rebuilt after deletion of sorted or random tuples, for trees that contain $n$ tuples. This table reveals that the largest subtree rebuilt for deletion of sorted tuples is 2 to 6 larger than the largest subtree rebuilt for deletion of random tuples. Rebuilding a larger subtree requires more time than rebuilding a smaller subtree.
@@ 1.7 AVL Versus Red-black Balance
Section 1.1 of this article explains that the balance of a k-d tree may be established either via the red-black criterion, which requires that the heights of the two child subtrees of a node differ by at most a factor of two; or via the AVL criterion, which requires that the heights of the two child subtrees of a node differ by 1. In fact, it is possible to define the AVL criterion less stringently; for example, to allow the subtrees to differ by an integer in the range $\left[1,4\right]$ instead of by 1 [Foster1973].
@@
Figure 8 plots the execution times (in microseconds) of search for the 1,000 nearest neighbors to tuple $\left(0,1,2\right)$ for trees that contain $n$ tuples. In this figure, the $x$-axis is $\log_{2}\left(n\right)$ because the worst-case performance of nearest-neighbor search is expected to be $O\left[\log_{2}\left(n\right)\right]$ [Friedman]. This figure reveals that the execution times for AVL balance exceed the execution time for red-black balance.
Figure 9 plots the execution times (in microseconds) of search for all tuples that lie within a cubic region that is centered at tuple $\left(0,1,2\right)$ and that occupies $1/1,000^{\mathrm{th}}$ of the 3-dimensional space, for trees that contain $n$ tuples. The $x$-axis in this figure is $n^{2/3}$ because the worst-case performance of region search is expected to be $O\left(n^{1-1/k}\right)=O\left(n^{2/3}\right)$ for $k=3$ [Lee]. This figure reveals that the execution times for AVL balance exceed the execution time for red-black balance.
@@
The improved search performance for red-black balance compared to AVL balance, as revealed by Figures 8 and 9, contradicts the expectation that search of an AVL tree should outperform search of an equivalent red-black tree because the height of the red-black tree is greater than the height of the AVL tree [Wirth]. And although Table 5 reveals that the heights of k-d trees balanced according to the red-black criterion exceed the heights of k-d trees balanced according AVL criteria for height differences of 1 to 4, red-black balance nevertheless outperforms AVL balance for nearest-neighbor search and region search.

### v14

@@ 1 Introduction
Bentley introduced the k-d tree as a binary search tree that stores k-dimensional data [2]. Like a binary search tree, a k-d tree partitions a set of data at each recursive level of the tree. Unlike a binary search tree that uses only one key for all levels of the tree, a k-d tree uses $k$ keys and cycles through the keys for successive levels of the tree. For example, to build a k-d tree from three-dimensional tuples that represent $\left(x,y,z\right)$ coordinates, the keys would be cycled as $x,y,z,x,y,z$... for successive levels of the k-d tree. A more elaborate scheme for cycling the keys chooses the coordinate that has the widest dispersion or largest variance to be the key for a particular level of recursion [8] but this article assumes that the keys are cycled as $x,y,z,x,y,z$... in order to simplify the exposition.
@@
Two dynamic k-d trees proposed previously maintain balance via a set of multiple k-d trees [15] or k-d* trees [19]. In contrast, the algorithms described in the present article maintain balance via a single k-d tree.
@@ 1.1 Two Definitions of Balance
This definition of height permits definition of two types of balance. The first type of balance, AVL balance, requires that each node’s left and right subtrees differ in height by at most one node [1] [19].
The second type of balance, red-black balance, requires that, at each node of the tree, the node’s left and right subtrees differ in height by a factor of 2 or less. This requirement derives from the following features of a red-black tree: (1) a node is colored either red or black, (2) neither child of a red node may be red, and (3) every path from any node to the bottom of the tree must contain the same number of black nodes [9]. Given these features, the shortest path from any node to the bottom of a red-black tree contains only black nodes, whereas the longest path contains alternating red and black nodes. Because every path from any node to the bottom of the tree must contain the same number of black nodes, the longest path is twice the length of the shortest path [5]. For a k-d tree, red-black balance doesn’t require coloring each node in the k-d tree via a color field. It requires only that each node’s left and right subtrees differ in height by a factor of 2 or less.
@@ 1.2 Insertion into a k-d Tree and Subsequent Rebalancing
Insertion into a k-d tree occurs in a similar manner to insertion into a standard binary search tree (i.e., a 1-d binary search tree) [5] [18]. To insert a k-dimensional tuple, the tree is searched recursively for that tuple, comparing the tuple’s super key to the tree node’s $x$:$y$:$z$, $y$:$z$:$x$, or $z$:$x$:$y$ super key at each level of the tree as the tuple descends the tree. If the search arrives at the bottom of the tree without finding the tuple, the tuple is inserted into the tree as a new leaf node.
Then, as the recursion unwinds, the height is computed at each node along the return path to the root of the tree, and the balance is checked at each node to determine whether rebalancing is required at that node. In Figure 1, the difference in child subtree heights is 0, 1, and 2 at nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ respectively. Hence, the subtree rooted at node $\left(8,1,5\right)$ must be rebalanced, as required by the AVL or red-black criterion because both criteria enforce a subtree height difference of at most one node. This subtree is rebalanced by rebuilding it [14] to create the subtree shown in Figure 2 that contains nodes $\left(9,2,1\right)$; $\left(8,3,2\right)$; and $\left(8,1,5\right)$ depicted with dashed borders, and whose respective $z$:$x$:$y$ super keys have relative order 11:99:22 << 22:88:33 << 55:88:11. Rebuilding a subtree that contains 3 nodes or fewer, such as this subtree, occurs frequently and requires comparison of at most 3 super keys [17]. Rebuilding a subtree that contains more than 3 nodes requires a static tree-building algorithm, such as a $O\left[n\log\left(n\right)\right]$ parallel algorithm described previously [3].
@@ 1.3 Deletion from a k-d Tree and Subsequent Rebalancing
Deletion from a k-d tree begins in a similar manner to, but is more complicated than, deletion from a 1-d binary search tree [5], [11] [18]. To delete a k-dimensional tuple, the tree is searched recursively for that tuple, comparing the tuple’s super key to the node’s $x$:$y$:$z$, $y$:$z$:$x$, or $z$:$x$:$y$ super key at each level of the tree. If the search finds the tuple, the tuple is deleted from the tree as one of three cases: (1) a leaf node with no children, (2) a node with one child, and (3) a node with two children. After deletion of the node, the height is computed at each node along the return path to the root of the tree as the recursion unwinds, and the balance is checked at each node to determine whether rebalancing is required at that node.
@@
The second case is deletion of a node that has one child. This case must be treated differently for a k-d tree than for a 1-d tree. In a 1-d tree, a 1-child node is replaced by its only child. However, for a k-d tree, replacing a 1-child node by its child violates the invariant of the subtree rooted at that child [5] [11]. A 1-child node, whose super key is $x$:$y$:$z$, has a child whose super key is $y$:$z$:$x$. So moving that child to the next higher level in the tree to replace its parent, as is done in a 1-d tree, would require that the child’s super key be $x$:$y$:$z$ instead of $y$:$z$:$x$.
This problem is solved by replacing a 1-child node by either its immediate predecessor node or its immediate successor node, similar to the replacement of a node that has two children in a 1-d tree. A 1-child node that has only a less-than child is replaced by its immediate predecessor, whereas a 1-child node that has only a greater-than child is replaced by its immediate successor [5] [11].
However, finding the predecessor or successor node is more complicated for a k-d tree than for a 1-d tree [5]. The predecessor node to a particular node in a 1-d tree is found by starting with the less-than child to that node, and recursively following the greater-than child pointer until it is null, because the null-pointer node has the largest key in the less-than subtree [20]. The successor node to a particular node is found in a similar manner by starting with the greater-than child to that node, and recursively following the less-than child pointer until it is null.
@@
Because node $b$ has a $y$:$z$:$x$ super key, its $x$:$y$:$z$ sorted order relative to node $a$ is unknown, so it must be inspected as a potential predecessor node to node $a$. Also, because of its $y$:$z$:$x$ super key, node $b$ cannot eliminate its children (nodes $c$ and $d$) from recursive search for node $a$’s predecessor node.
@@
In view of the exclusion of less-than branches, finding 1-child node $a$’s immediate predecessor node requires inspecting $n^{1-1/k}$ nodes of the $n$-node subtree rooted at node $b$, assuming that subtree to be perfectly balanced with no empty branches [16]. And if 1-child node $a$ has only a greater-than child instead of a less-than child, finding node $a$’s immediate successor node requires searching the subtree rooted at that greater-than child to find the node that has the smallest $x$:$y$:$z$ super key.
When the immediate predecessor or immediate successor node to the 1-child node has been found in the manner described above, that predecessor or successor node’s tuple field is copied to the 1-child node’s tuple field, and then the predecessor or successor node is deleted from the tree. But that predecessor or successor node is not necessarily a leaf node, and if not, it must be deleted recursively as one of the three cases itemized above [11], [16]. Recursive deletion converges to a leaf node that is excised from the tree. Then the recursion unwinds, with computation of the height and potential rebalancing at each node along the return path.
The third case is deletion of a node that has two children. This case differs from the 1-child case only in that a deleted 2-child node may be replaced by either its immediate predecessor or its immediate successor node. The choice between the predecessor or successor is arbitrary. However, the optimum replacement node might be selected from the 2-child node’s child subtree that has the greater height, in an attempt to improve balance and thereby minimize rebalancing the subtree rooted at the 2-child node [6].
@@ 1.4 Benchmark Methodology
A k-d tree benchmark measured algorithm performance for trees that contained $n$ nodes, for values of $n$ in the range $\left[1,003,201;4,523,071\right]$ that map to equally spaced values of $n\log_{2}\left(n\right)$ in the range $\left[20,000,000;100,000,000\right]$. Each node of the tree stored a $k$-dimensional tuple of 64-bit integers. The integers were equally spaced across the maximum 64-bit integer range $r$, so the spacing was $r/n$. The integers were randomly shuffled via the std::mt19937_64 Mersenne Twister pseudo-random number generator [13] and copied to the first of the $k$ dimensions, then randomly shuffled again and copied to the second of the $k$ dimensions, et cetera. So that all benchmarks randomly shuffled the integers in an identical sequence, each benchmark initialized std::mt19937_64 to std::mt19937_64::default_seed.
@@
Each benchmark measured the execution times for insertion of all tuples into the tree, verification of correct ordering of the tree, search for all tuples in the tree, and deletion of all tuples from the tree via the std::chrono::steady_clock::now() function. To provide a comparison, the execution time was measured for construction of a static k-d tree via a $O\left[n\log\left(n\right)\right]$ algorithm [3]. Each benchmark was repeated 100 times and the mean values and standard deviations of the execution times were calculated. The standard deviations were less than 1% of the mean values.
To assess the performance of the dynamic k-d tree for worst-case data, a set of tuples was created by constructing a static k-d tree, and then sweeping that tree in sorted order to obtain a set of tuples to insert into and delete from a dynamic k-d tree.
@@ 1.5 Benchmark Results
In Figure 4, the execution-time plots for insertion, deletion, and search of random and sorted tuples in a dynamic k-d tree, and for creation of a static k-d tree from random tuples, exhibit $O\left[n\log\left(n\right)\right]$ computational complexity amortized over $n$ tuples, as expected for a balanced binary tree. The execution time for insertion of random tuples is about 1.5 times longer than the execution time for creating a static k-d tree en masse from random tuples. This disparity represents the difference between rebalancing piecemeal for a dynamic tree, versus balancing en masse for a static tree.
Figure 4 reveals that the execution time for insertion of random tuples exceeds the execution time for insertion of sorted tuples. This difference in execution times is likely due to cache memory effects. For insertion of sorted 3-dimensional $\left(x,y,z\right)$ tuples, two consecutive tuples descend the nascent tree via a similar path. Hence, the nodes that the first tuple causes to be loaded into cache during its descent of the tree likely remain resident in cache and are likely accessed by the second tuple during its descent of the tree. In contrast, for insertion of random tuples, it is unlikely that two consecutive tuples descend the tree via a similar path. Hence, the nodes loaded into cache during the first tuple’s descent of the tree are unlikely to be accessed by the second tuple. A similar disparity in execution times between sorted and random keys has been observed for the AVL and red-black trees [4].
Figure 4 also reveals that the execution time for deletion of sorted tuples exceeds the execution time for deletion of random tuples. Compared to insertion of sorted tuples, it is less likely for deletion of sorted tuples that two consecutive tuples descend the tree via a similar path. This lower probability of a similar path arises because, during the prior insertion of sorted tuples, the tree was considerably restructured via rebalancing. So it is unlikely that, during deletion, nodes loaded into cache during the first tuple’s descent of the tree will be accessed by the second tuple. Hence, it is not expected that deletion of sorted tuples would be faster than deletion of random tuples.
Moreover, Table 1 provides a possible reason that deletion of sorted tuples is slower than deletion of random tuples. This table reveals that the largest subtree rebuilt for deletion of sorted tuples is 2 to 6 times larger than the largest subtree rebuilt for deletion of random tuples. Rebuilding a larger subtree requires more time than rebuilding a smaller subtree.
@@ 1.7 AVL Versus Red-black Balance
Section 1.1 of this article explains that the balance of a k-d tree may be established either via the red-black criterion, which requires that the heights of the two child subtrees of a node differ by at most a factor of two; or via the AVL criterion, which requires that the heights of the two child subtrees of a node differ by 1. In fact, it is possible to define the AVL criterion less stringently; for example, to allow the subtrees to differ by an integer in the range $\left[1,4\right]$ instead of by 1 [7].
@@
Figure 8 plots the execution times (in microseconds) of search for the 1,000 nearest neighbors to tuple $\left(0,1,2\right)$ for trees that contain $n$ tuples. In this figure, the $x$-axis is $\log_{2}\left(n\right)$ because the worst-case performance of nearest-neighbor search is expected to be $O\left[\log_{2}\left(n\right)\right]$ [8]. This figure reveals that the execution times for AVL balance exceed the execution time for red-black balance.
Figure 9 plots the execution times (in microseconds) of search for all tuples that lie within a cubic region that is centered at tuple $\left(0,1,2\right)$ and that occupies $1/1,000^{\mathrm{th}}$ of the 3-dimensional space, for trees that contain $n$ tuples. The $x$-axis in this figure is $n^{2/3}$ because the worst-case performance of region search is expected to be $O\left(n^{1-1/k}\right)=O\left(n^{2/3}\right)$ for $k=3$ [12]. This figure reveals that the execution times for AVL balance exceed the execution time for red-black balance.
@@
The improved search performance for red-black balance compared to AVL balance, as revealed by Figures 8 and 9, appears to contradict the expectation that search of an AVL tree should outperform search of an equivalent red-black tree because the average height of a red-black tree is greater than the average height of the equivalent AVL tree [20]. And although Table 5 reveals that the heights of k-d trees balanced via the red-black criterion exceed the heights of k-d trees balanced via the AVL criteria for height differences of 1 to 4, red-black balance nevertheless outperforms AVL balance for nearest-neighbor search and region search. A possible explanation for this discrepancy is that Table 5 reports maximum heights, whereas the expectation of AVL tree outperformance is based on a consideration of average heights.